OpenStudy (anonymous):

If the product of two consecutive integers is decreased by twenty times the greater integer, the result is 442. Find the integers.

OpenStudy (cwrw238):

let integers be n and n+1 writing it as an equation: n(n+1) - 20(n+1) = 442 n^2 + n - 20n - 20 = 442 n^2 - 19n - 462 = 0 solve for n

OpenStudy (cwrw238):

this might factor: 462 = 2*231 = 2*7*3*11

OpenStudy (anonymous):

so 14 or 33

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

yup - -19 = -33 + 14 so what is n?

OpenStudy (anonymous):

14 or 33..?

OpenStudy (anonymous):

Isn't that n?

OpenStudy (anonymous):

*-14

OpenStudy (cwrw238):

the factors are (n- 33)(n + 14) = 0 so n is 33 or -14

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

yw